# Area Law for the entanglement entropy of free fermions in nonrandom ergodic field

### v2

@@ 3. Finite-difference operators with quasi-periodic potential
Now the triangle inequality $|\mathbf{m-l}|+|\mathbf{n-l}|\geq|\mathbf{m-n}|$ yields the following upper bound for the r.h.s.
@@
Remark 3.1.
The description (3.1) – (3.3) of eigenfunctions of the Maryland model provides the simplest and a quite explicit illustration of the notion of the localisation centers and the uniformly localised eigenfunctions. These spectral objects can be defined for any selfadjoint operator $H$ acting on $\ell^{2}(\mathbb{Z}^{d})$. Namely, $H$ is said to have uniformly localised eigenfunctions ((ULE)) if and only if $H$ has a complete set
Remark 3.2.
@@
  • the whole spectrum of $H_{\omega}$ is $\{\lambda_{\mathbf{l}}(\omega)\}_{\mathbf{l}\in\mathbb{Z}^{d}}$, where $\lambda_{\mathbf{l}}(\omega)=\lambda(\omega+\langle\alpha,\mathbf{l}\rangle)$, and the corresponding eigenfunctions are
    (3.14) $\psi_{\mathbf{l}}(\omega)=\{\psi_{\mathbf{l}}(\omega,\mathbf{n})\}_{\mathbf{n}\in\mathbb{Z}^{d}},\;\psi_{\mathbf{l}}(\omega\mathbf{,n})=\psi(\omega+\langle\alpha,\mathbf{l}\rangle;\mathbf{n-\mathbf{l}}),$
@@
Thus, the description of the spectrum for this class of operators coincides with that of (3.1) – (3.3) for the Maryland model and admits the exponential (ULE) bound (3.11) analogous to (3.4). Thus, repeating literally the proof of Theorem 1 (i), we arrive to the same conclusion on the validity of the Area Law (and the exponential dynamical localisation in expectation) on the whole spectrum of the operators in question. ∎
@@
The spectrum of the corresponding operator (the one-dimensional Schrödinger operator (2.12) with potential (2.14)) is also described by formulae (3.1) – (3.3). However, since in this case the operator $W$ in (2.1) – (2.2) is the one-dimensional discrete Laplacian, an important part of the proof of (3.4) can be made explicit, thereby leading to its following one-dimensional version
with $\omega$-independent parameters $C\leq\infty,\;c>0$ and $\varepsilon_{0}\geq 0$, see Appendix B for the proof of this bound. This allows us to apply again formulae (3.6) – (3.9) to obtain (2.22) and then Criterion 1 implies the assertion of Theorem 2 (i). ∎
@@
This bound is similar to the one of [28], but is valid for all $\omega$. However, unlike (3.4), or, more generally, (3.5), the bound (3.16) cannot be used immediately to prove the exponential decay of the Fermi projector and/or the eigenfunction correlator, as was possible with the estimate (3.4) and(3.5), see formulae (3.4) – (3.9). The reason is that the functions $\{u_{l}\}_{l}$, although orthogonal, are not orthonormalized. Therefore, an additional argument must be used to pass from $\{u_{l}\}_{l}$ to $\{\psi_{l}\}_{l}$ in the above bound. Such an elegant argument was proposed in [28]. Thus, by using this argument and (3.16), it is possible to prove the exponential dynamical localisation in expectation as well as the exponential decay of the Fermi projection, the eigenfunction correlator, and the assertion of Theorem 2 $(ii)$. ∎
@@ 4. Schrödinger operators with potentials generated by hyperbolic transformations
4.1. Preliminaries
@@ 4.1.2. Assumptions and definitions
4.2. Proof of Theorems 3 and 4
To prove the theorems we start with proving several assertions.
The first assertion (Lemma (4.4)) provides a useful criterion for the exponential decay of the eigenfunction correlator corresponding to a bounded ergodic one-dimensional discrete Schrödinger operator. We believe that the criterion is of independent interest.
Let $M=\{M(j,k)\}_{j,k=-n}^{n}$ be a $(2n+1)\times(2n+1)$ matrix. Denote by
where $(M-\lambda)^{-1}(k,l)$ is the $(k,l)$ element of the matrix $(M-\lambda)^{-1}$, the set of the “bad” spectral parameters.
Lemma 4.4.
$Q_{I}(0,l)\leq 16e^{-\frac{c}{20}l}\equiv 16e^{-c^{\prime}l}.$
Proof.
Let $G$ be the resolvent operator of $H$ and $G_{[m,p]}$ be the resolvent operator of the restricted operator $H^{[m,p]}$. Let $n$ be such that $n=l/20$ and assume that $k=0$. The second resolvent identity implies for each $i,j\in\{0,1\}$
and
$|G(0,l)|\leq e^{-cn}\big(|G(n-j+1,l)|+|G(-n+i-1,l)|\big),$
$|G(0,l)|\leq e^{-cn}\big(|G(0,l-n+i-1)|+|G(0,l+n-j+1)|\big).$
Recall that [3]
When we take $\epsilon\to 0$ and $\Lambda\nearrow\mathbb{Z}$ we obtain
where in the second inequality we used the fact that $Q_{I}(n,m)\leq 1$ for any $n,m\in\mathbb{Z}$ and the last one follows from the choice of $n$. ∎
Next, we have a large deviation type estimate for the Schrödinger operators with potentials generated by the subshift of finite type under the assumptions of Theorems 3 and 4.
@@ 4.2. Proof of Theorems 3 and 4
Now we are ready to prove Theorems 3 and 4.
Proof of Theorems 3 and 4.
@@
As we indicated in the preliminaries, there is a one-to-one correspondence between subshifts of finite type and Markov chains. To prove (4.24), we will pass to the corresponding Markov chain setting. Let $S=\{1,\dots,k\}$ be the phase space of the Markov chain. Since $\mathbf{P}$ is $T$-ergodic, we conclude that the corresponding transition matrix $P$ is irreducible, hence, the Markov chain is irreducible. Since $T$ has a fixed point then $P$ and thus the Markov chain are aperiodic, namely there exists an integer $m$ such that for all $1\leq i\leq k$ we have $(P^{m})_{ii}>0$. Thus, (4.24) follows from the Convergence Theorem (see e.g. [36, Theorem 4.9]).
@@
where the second inequality follows from Lemma 4.4 and (4.25).
Thus, by Criterion 2 the corresponding entanglement entropy in this case obeys the Area Law at the bottom of the spectrum. ∎
@@
Appendices
Appendix A Proof of Bound (3.4).
Note first that here the function $\lambda_{0}$ of (3.2) is the functional inverse of the Integrated Density of States $N(\lambda)$ of $H$ (see [38, Sections 4.B – 4.C] and [3, Sections 3.3 – 3.4] for its definition and properties). We have for the Maryland model
(A.1) $N(\lambda)=\int_{-\infty}^{\lambda}n(\lambda)d\lambda,$
i.e.,
(A.2) $\lambda_{0}\circ N=1,$
where
(A.3) $n(\lambda)=\frac{g}{(2\pi i)^{d}\pi}\int_{\mathbb{T}^{d}}|w(\mathbf{\eta})-\lambda+ig|^{-2}\prod_{j=1}^{d}\frac{d\eta_{j}}{\eta_{j}},$
is the Density of States of $H$ and $w:\mathbb{T}^{d}\rightarrow\mathbb{R}$ is given by the $d$-dimensional Fourier series with coefficients of (2.2)
(A.4) $\mathbf{\eta}$ $=(\eta_{1},\dots,\eta_{d})\in\mathbb{T}^{d},\,\,\mathbf{n}=(n_{1},\dots,n_{d})\in\mathbb{Z}^{d}.$
Next, the function $\psi_{0}$ of (3.1) is (see [38, Section 18.B])
(A.5) $\psi_{0}(\lambda,{\mathbf{n}})=\frac{1}{(2\pi i)^{d}(\pi n(\lambda)/g)^{1/2}}\int_{\mathbb{T}^{d}}\frac{e^{t(\lambda,\mathbf{\eta})}}{(w(\mathbf{\eta})-\lambda-ig)}\prod_{j=1}^{d}\eta^{n_{j}-1}d\eta_{j},$
where
(A.6) $t(\lambda,\mathbf{\eta})=\sum_{\mathbf{m}\in\mathbb{Z}^{d}}t_{\mathbf{m}}(\lambda)\prod_{j=1}^{d}\eta_{j}^{m_{j}},\;t_{\mathbf{m}}(\lambda)=\frac{l_{\mathbf{m}}(\lambda)}{(1-e^{i\langle\mathbf{\alpha,m}\rangle})}$
with
(A.7) $l_{\mathbf{m}}(\lambda)=(2\pi i)^{-d}\int_{\mathbb{T}^{d}}\log c(\lambda,\mathbf{\eta})\prod_{j=1}^{d}\eta^{m_{j}-1}d\eta_{j},$
and
(A.8) $c(\lambda,\mathbf{\eta})=-\frac{w(\mathbf{\eta})-\lambda+ig}{w(\mathbf{\eta})-\lambda-ig}.$
Thus, the $c(\lambda,\mathbf{\eta})$ of (A.8) is the input of the above analytic procedure given by formulae from (A.7) to (A.5) that leads to the eigenfunctions (3.3).
Our goal is to prove the bound (3.4) by using these formulae, that allows us, using Remark 3.2, to apply Criterion 2 to establish the validity of the Area Law.
We are going to use the following simple fact.
Lemma A.1.
Let $f:\mathbb{T}^{d}\rightarrow\mathbb{C}$ be a periodic function and $\{F_{\mathbf{n}}\}_{\mathbf{n}\in\mathbb{Z}^{d}}$ be its Fourier coefficients.
where for any $0<\rho_{2}<\rho_{1}$ and $a=\rho_{1}-\rho_{2}>0$ it has the bound
[…]
(A.10) $F$ $=$ $\max_{|r|\leq\rho_{1},\;|\mathbf{\eta}|=1}|f(e^{r}\mathbf{\eta})|.$
Proof.
Now, applying Lemma A.1 (i) to $w$ of (A.4) whose Fourier coefficients satisfy (2.2), we find that $w$ admits the analytic continuation to $\mathbb{T}_{\rho_{1}}^{d},\;0<\rho_{1}<\rho$, where $\rho>0$ is given in (2.2). Moreover, since the “initial” $w$ of (A.4) is real-valued ($w:\mathbb{T}^{d}\rightarrow\mathbb{R}$), we have for its analytic continuation
(A.11) $\lim_{r\rightarrow 0}\mathrm{Im}\ w(e^{r}\mathbf{\eta})=0,\;|\mathbf{\eta}|=1.$
(A.12) $|w(e^{r}\mathbf{\eta})-\lambda\pm ig|\geq|g\pm\mathrm{Im}\,w(e^{r}\mathbf{\eta})|\geq|g|-|\mathrm{Im}\,w(e^{r}\mathbf{\eta})|\geq g/2,$
where the last inequality follows from (A.11). We have then from (A.8) that $|\log c(\lambda,\mathbf{\eta})|\leq|\log|c(\lambda,\mathbf{\eta})|\ |+2\pi,$ and for any $|r|\leq\rho_{2}$
$\leq$ $2|g|\ |w(e^{r}\mathbf{\eta})-\lambda-ig|^{-1}\leq 4,$
where in obtaining the last bound of the r.h.s. above we used (A.12). Combining the two last bounds, we get for any $|r|\leq\rho_{2}$
$|\log c(\lambda,e^{r}\mathbf{\eta})|\leq C,$
where $C$ is a constant. Combining (A.7) and Lemma A.1 (ii) we obtain
$|l_{\mathbf{m}}(\lambda)|\leq Le^{-\rho_{2}|\mathbf{m}|},$
where $L<\infty$ and $\rho_{2}>0$ do not depend on $\lambda$ and $\mathbf{m}$.
This and the Diophantine condition (2.5) implies for the Fourier coefficients $t_{\mathbf{m}}(\lambda)$ of (A.6)
(A.13) $|t_{\mathbf{m}}(\lambda)|\leq Le^{-\rho_{2}^{\prime}|\mathbf{m}|},\;0<\rho^{\prime}_{2}<\rho_{2}.$
Then, Lemma A.1 (ii) imply that $t(\lambda,\cdot)$ of (A.6) admits the analytic continuation into $\mathbb{T}_{\rho_{3}}^{d}$ with some $0<\rho_{3}<\rho^{\prime}_{2}$ and is bounded there for any $|r|\leq\rho_{3}$, i.e.,
(A.14) $|t(\lambda,e^{r}\mathbf{\eta})|\leq T<\infty,\;|r|\leq\rho_{3},$
where $T$ and $\rho_{3}$ are $\lambda$-independent according to (A.9).
Next, (A.12), and the analyticity of $w$ in $\mathbb{T}_{\rho_{1}}^{d}$ imply that
is analytic in $\mathbf{z}$ in $\mathbb{T}_{\rho_{3}}^{d}$ for any $\lambda\in\mathbb{R}$. Since the r.h.s. of (A.5) is the Fourier coefficient of $h(\lambda,\mathbf{\cdot})$, (A.10) yields the following bound for the r.h.s. of (A.5)
(A.15) $\Psi e^{-\rho_{3}|\mathbf{n}|},\;\mathbf{n}\in\mathbb{Z}^{d},\;\Psi=\max_{\lambda\in\mathbb{R},\;|r|\leq\rho_{3},\;|\mathbf{\eta}|=1}|h(\lambda,e^{r}\mathbf{\eta})|.$
It follows from (A.3) that
  1. (1)
    $(n(\lambda))^{-1/2}\geq(\pi g)^{1/2},$ and $(n(\lambda))^{1/2}=|\lambda|^{-1}(1+o(1)),\;|\lambda|\rightarrow\infty$, by (A.3);
  2. (2)
    $|w(\mathbf{z})-\lambda-ig|,\;|\mathbf{z}|\in[e^{-\rho_{3}},e^{\rho_{3}}]$ is bounded from below by (A.12) and is $|\lambda|(1+o(1)),\;|\lambda|\rightarrow\infty$;
  3. (3)
    $e^{t(\lambda,\mathbf{z})}\leq e^{T}$ by (A.14).
This and (A.15) imply that $\Psi$ is independent of $\lambda\in\mathbb{R}$ and $\mathbf{n}\in\mathbb{Z}^{d}$.
We obtained (3.4) with $\Psi$ of (A.15) as $C$ and $\rho_{3}$ of (A.14) as $c$.
Appendix B Proof of Bound (3.15).
We will use here several basic facts from [38, Section 18.C].
Since in this case operator $W$ in (2.1) is the one-dimensional discrete Laplacian, we set $W_{0}=0,\;W_{\pm 1}=1$ in (2.2). Hence, $w$ of (A.4) is
(B.1) $w(\eta)=\eta+\eta^{-1}.$
Thus, $w$ is analytic everywhere except zero and infinity, i.e., in this case the analog of $\rho_{1}$ of the previous appendix is infinity.
Next, the integral in the r.h.s. of formula (A.7) for $d=1$ and $w$ of (B.1) can be calculated yielding (cf. (A.7))
$l_{m}=\frac{2i}{|m|}e^{-\gamma|m|}\sin|m|\varphi.$
(B.2) $u_{n+1}+u_{n-1}+g\tan(\alpha n+\omega)u_{n}=\lambda u_{n}.$
We have
(B.3) $\mathrm{sinh}\gamma(\lambda,g)=s>0,$
where $s>0$ is such that
$s^{4}+(4-\lambda^{2}-g^{2})s^{2}-g^{2}=0,$
$\gamma(\lambda,g)\geq\gamma(0,g)=\mathrm{arc\,sinh}(|g|/2)>0,\;g\neq 0.$
$\rho^{\prime}:=\gamma(\varepsilon_{0},g)-\beta(\alpha)>0,$
we obtain from (A.6) and (A.7) for the Fourier coefficients $t_{m}(\lambda)$ of function $t(\mathbf{\eta},\cdot)$ of (A.6) for $d=1$:
This bound is an analog of (A.13). Hence, repeating the argument of the previous appendix starting from (A.6) we get (3.15).
Appendix C Another Proof of Bound (2.22) for the Maryland model.
@@
Next, it follows from (A.1) – (A.2) that $d\omega=n(\lambda)d\lambda$, hence, the r.h.s. above is
@@
In particular, using the indicator $\chi_{(-\infty,\varepsilon_{F}]}$ as $\phi$, we obtain having (A.1) (cf. (3.4))

### v3

@@ 3.2. Proof of Theorem 1
Now the triangle inequality $|\mathbf{m-l}|+|\mathbf{n-l}|\geq|\mathbf{m-n}|$ yields the following upper bound for the r.h.s. (3.28)
@@
  • the spectrum of $H_{\omega}$ occupies the whole $\mathbb{R}$, is pure point and of multiplicity 1, its eigenvalues $\{\lambda_{\mathbf{l}}(\omega)\}_{\mathbf{l}\in\mathbb{Z}^{d}}$ and the corresponding eigenfunctions $\{\psi_{\mathbf{l}}(\omega)\}_{\mathbf{l}\in\mathbb{Z}^{d}}$, $\psi_{\mathbf{l}}(\omega)=\{\psi_{\mathbf{l}}(\omega,\mathbf{n})\}_{\mathbf{n}\in\mathbb{Z}^{d}}$ are
    (3.32) $\lambda_{\mathbf{l}}(\omega)=\lambda(\omega+\langle\mathbf{\alpha},\mathbf{l}\rangle),\;\psi_{\mathbf{l}}(\omega,\mathbf{n})=\psi(\omega+\langle\mathbf{\alpha},\mathbf{l}\rangle,\mathbf{n}-\mathbf{l}),$
@@
We conclude that the description of the spectrum for this class of operators coincides with that of (3.1) – (3.3) for the Maryland model and admits the exponential (ULE) bound (2.23) analogous to (2.20). Thus, repeating literally the proof of Theorem 1 (i), we arrive to the same conclusion on the validity of the Area Law (and the exponential dynamical localisation in expectation) on the whole spectrum of the operators in question. ∎
3.3. Proof of Theorem 2
@@
The estimate (3.26) allows us to apply again formulae (3.27) – (3.30) to obtain (2.29) and then Criterion 1 implies the assertion of Theorem 2 (i). ∎
@@
This bound is similar to the one of [29], but is valid for all $\omega$. However, unlike (3.26), or, more generally, (3.20), the bound (3.33) cannot be used immediately to prove the exponential decay of the Fermi projector and/or the eigenfunction correlator, as was possible with the estimate (2.20) and(3.20), see formulae (2.20), (3.30). The reason is that the functions $\{u_{l}\}_{l}$, although orthogonal, are not orthonormalized. Therefore, an additional argument must be used to pass from $\{u_{l}\}_{l}$ to $\{\psi_{l}\}_{l}$ in the above bound. Such an elegant argument was proposed in [29]. Thus, by using this argument and (3.33), it is possible to prove the exponential dynamical localisation in expectation as well as the exponential decay of the Fermi projection, the eigenfunction correlator, and the assertion of Theorem 2 $(ii)$. ∎
@@
4. Proof of general criterion for “strong” dynamical localisation
Here we prove Lemma 2.5. Let $G$ be the resolvent operator of $H$ and $G_{[m,p]}$ be the resolvent operator of the restricted operator $H^{[m,p]}$. Let $n$ be such that $n=l/20$ and assume that $k=0$. The second resolvent identity implies for each $i,j\in\{0,1\}$
and
$|G(0,l)|\leq e^{-cn}\big(|G(n-j+1,l)|+|G(-n+i-1,l)|\big),$
$|G(0,l)|\leq e^{-cn}\big(|G(0,l-n+i-1)|+|G(0,l+n-j+1)|\big).$
Recall that [3]
When we take $\epsilon\to 0$ and $\Lambda\nearrow\mathbb{Z}$ we obtain
where in the second inequality we used the fact that $Q_{I}(n,m)\leq 1$ for any $n,m\in\mathbb{Z}$ and the last one follows from the choice of $n$.
@@ 5. Schrödinger operators with potentials generated by hyperbolic transformations
5.1. Necessary facts of the subshifts of finite type (SFT) and the SFT-driven Schrödinger operators
@@ 5.1.2. Assumptions and definitions
5.2. Proof of Theorems 4 and 5
We start with several necessary assertions.
First, we prove the following large deviation type estimate for the Schrödinger operators with potentials generated by the subshift of finite type under the assumptions of Theorems 4 and 5.
@@ 5.2. Proof of Theorems 4 and 5
Now we are ready to prove Theorems 4 and 5.
Proof of Theorem 4.
@@
As we indicated in the preliminaries, there is a one-to-one correspondence between subshifts of finite type and Markov chains. To prove (5.23), we will pass to the corresponding Markov chain setting. Let $S=\{1,\dots,k\}$ be the phase space of the Markov chain. Since $\mathbf{P}$ is $T$-ergodic, we conclude that the corresponding transition matrix $P$ is irreducible, hence, the Markov chain is irreducible. Since $T$ has a fixed point then $P$ and thus the Markov chain are aperiodic, namely there exists an integer $m$ such that for all $1\leq i\leq k$ we have $(P^{m})_{ii}>0$. Thus, (5.23) follows from the Convergence Theorem (see e.g. [39, Theorem 4.9]).
@@
where the second inequality follows from Lemma 2.5 and (5.24). ∎
Proof of Theorem 5.
It follows from Theorem 4 and Criterion 2 that the corresponding entanglement entropy in this case obeys the Area Law at the bottom of the spectrum. ∎
@@
Appendix A Another Proof of Bound (2.29) for the Maryland model.
@@
Next, it follows from (3.4) – (3.5) that $d\omega=n(\lambda)d\lambda$, hence, the r.h.s. above is
@@
In particular, using the indicator $\chi_{(-\infty,\varepsilon_{F}]}$ as $\phi$, we obtain having (3.4) (cf. (2.20))